# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ B-C Schrodinger’s Equation
B-D Hermitian Matrix/Operator
@@ B-D1 Hamiltonian
B-E Observable
Projective measurements are a special case of general measurements that have the ability to perform unitary transformations. A projective measurement is described by an observable (M) and a Hermitian operator on the state space of the system being observed. The observable has a spectral decomposition. The completeness relation, whose probabilities sum to 1, is applicable to projective measurements and makes the probability constraint valid. This way of computing measurements is related to the Heisenberg uncertainty principle [36].
B-F Heisenberg’s Uncertainty Principle

### v3

@@ B.3 Schrodinger’s Equation
B.4 Hermitian Matrix/Operator
@@ B.4.1 Hamiltonian
B.5 Observable
Projective measurements are a special case of general measurements that have the ability to perform unitary transformations. A projective measurement is described by an observable (M) and a Hermitian operator on the state space of the system being observed. The observable has a spectral decomposition. The completeness relation, whose probabilities sum to 1, is applicable to projective measurements and makes the probability constraint valid. This way of computing measurements is related to the Heisenberg uncertainty principle (Nielsen & Chuang, 2010b).
B.6 Heisenberg’s Uncertainty Principle